OpenStudy (anonymous):

Find the Derivative, Using Chain Rule. g(x)=9/5x^2

zepdrix (zepdrix):

Hey Phi!\[\Large\rm g(x)=\frac{9}{5x^2}\]Is this what the function looks like? With the x^2 in the bottom?

OpenStudy (anonymous):

Yes :D

zepdrix (zepdrix):

Changing some things up before finding a derivative, \[\Large\rm =\frac{9x^{-2}}{5}=\frac{9}{5}x^{-2}\]Ok with those steps? This form allows us to easily apply our power rule.

zepdrix (zepdrix):

Imma go make some foods :3 You think about it!

OpenStudy (anonymous):

Okay, so i know you cant have an exponet in the demoninator which is why you brought it up. then you should multiply the -2 to the constant so it would be \[\frac{ -18 }{ 5 }x^-3 \] ?

OpenStudy (nincompoop):

so how is that a chain rule?

OpenStudy (anonymous):

That what i dont get, it says apply chain rule. i dont know where.

OpenStudy (nincompoop):

I think you're missing an extra step, which does not come intuitively. It is something you need to be creative about

OpenStudy (anonymous):

I dont even know.

OpenStudy (nincompoop):

what does chain rule look like? you need to make your original function look like something you can apply chain rule to

OpenStudy (anonymous):

but it doesnt look like the chain rule would work there.

OpenStudy (nincompoop):

\(\huge g(x)=\frac{9}{5x^2} \rightarrow \frac{9}{\color{red}{\sqrt{25x^4}}}\)

OpenStudy (nincompoop):

does that look like something you can apply chain-rule to now?

OpenStudy (anonymous):

how did you get the demoninator? because i was taught that a Sqrt musnt be in the bottom?

OpenStudy (nincompoop):

it is not that SQUARE ROOT MUST NOT BE in the BOTTOM, it is just mathematically preferred for it in the numerator, I didn't really change anything. The value is still the same, I just looked for ways a chain-rule can be applied by first understanding what a chain-rule is and then make my problem looking similar to it without changing its value.

OpenStudy (nincompoop):

in calculus, you will see a lot of square roots in the denominator

OpenStudy (anonymous):

Okay, understood. Now you must use quotient rule correct? then apply chain rule? or is that wrong?

OpenStudy (nincompoop):

you're correct in your quotient rule, \(\large g(x) = \frac{u}{v} \rightarrow \huge g'(x) = \frac{u'v-uv'}{v^2} \) the derivative of your denominator uses chain-rule.

OpenStudy (anonymous):

How would you find the derivative of \[\sqrt{25x^4}\]

OpenStudy (nincompoop):

use chain rule

OpenStudy (nincompoop):

LAUGHING OUT LOUD!

OpenStudy (anonymous):

Lol, how? i dont understand

OpenStudy (nincompoop):

\(u=\sqrt{25 x^4} \rightarrow \huge (25x^4)^\frac{1}{2} \) chain-rule states that we take the derivative of the outer function \( \huge\frac{1}{2}(25x^4)^{-\frac{1}{2}} \) then multiply it by the derivative of the inner function \(\huge 4\times 25x^3 \) giving us \( \color {blue}{\huge u' = \frac{1}{2}(25x^4)^{-\frac{1}{2}} \times 100x^3} \) now just simplify so it looks nice and clean

OpenStudy (nincompoop):

I thought you knew chain-rule already

OpenStudy (nincompoop):

I used u instead of v... just switch it around

OpenStudy (anonymous):

In a simple problem i find it easy, but using it with quotient is harder for me

OpenStudy (nincompoop):

you can turn a quotient into a product

OpenStudy (anonymous):

but i thought quotient need a demoninator?

OpenStudy (nincompoop):

ye g(x) = u/v u is the numerator v is the denominator so you have two separate functions that makes up the quotient, the u and the v

OpenStudy (anonymous):

im lost.

OpenStudy (nincompoop):

I don't know why have you learned quotient rule?

OpenStudy (anonymous):

Recently yes.

OpenStudy (nincompoop):

so what is the problem and why are you lost?

OpenStudy (nincompoop):

do you remember what it is?

OpenStudy (anonymous):

the way you did it above for the demoninator made is complex, and i dont understand what was happening/

OpenStudy (nincompoop):

your problem states to solve using chain rule the way how it was originally written, chain rule cannot be used so I made it look like something chain rule can be used without altering its value it looks different now, but the value is still the same it is like saying \(2 \times 2 = 2^2 \) they are written differently, but the value is the same I just wrote the original problem into a format I can use chain rule

OpenStudy (nincompoop):

if you can't do fractional exponents and if algebra is not your strong suit, you will have a lot of trouble in calculus.

OpenStudy (nincompoop):

I've put a lot of time explaining already. It is time to move on.

OpenStudy (anonymous):

I understand. Ugh a junior shouldnt be taking Ap-Calculus, it's very hard. Thanks for helping though.

OpenStudy (nincompoop):

you're a junior highschool, you should be taking calculus at that level like how other kids in other countries do.

OpenStudy (nincompoop):

you just need to sharpen up and keep practicing go back to algebra or pre-calculus and make sure you master everything I bet you that once you've achieved it, you won't have trouble in calculus

OpenStudy (nincompoop):

good luck cheers

OpenStudy (anonymous):

Thank You